Use a triple integral to determine the volume of the region bounded by z =


p


x


2 + y


2 and z = x


2 + y


2


In 1st octant


1
Expert's answer
2021-09-20T18:21:20-0400

To convert from cylindrical to rectangular coordinates, we use the equations


"x=r\cos \theta, y=r\sin \theta, z=z"

"z=\sqrt{x^2+y^2}, z=\sqrt{(r\cos \theta)^2+(r\sin \theta)^2}=r"

"z=x^2+y^2, z=(r\cos \theta)^2+(r\sin \theta)^2=r^2"

"r=r^2, r_1=0, r_2=1"

"V=\displaystyle\int_{0}^{\pi/2}d\theta\displaystyle\int_{0}^{1}rdr\displaystyle\int_{r^2}^{r}dz"

"=\displaystyle\int_{0}^{\pi/2}d\theta\displaystyle\int_{0}^{1}(r^2-r^3)dr"

"=\dfrac{\pi}{2}(\dfrac{1}{3}-\dfrac{1}{4})=\dfrac{\pi}{24}({units}^2)"